# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust in Healthcare

### v2

@@ 4 Results
Study 1 was completed by 99 participants, from which 1 reported to have colour-blindness (we did not measure demographics). Since 5 participants failed the attention check, we did not include their data in any of our analyses (see more details below) resulting in a final sample of 94. Study 2 was completed by 69 medical students, from which 3 reported to have colour-blindness. [^8] One participant skipped the survey in Study 1 and one participant left a non-university email in Study 2; we deleted the data of both participants and did not include them in the results. The mean age of the participants in Study 2 was 22.3 (SD = 3.3), with 35 females (32 males, 1 other and 1 preferred not to disclose). As expected, participants in the medical student sample (Study 2) reported to have heard more about medical algorithms than participants in the Prolific sample (Study 1), but the number of people who had tried a medical AI tool before was negligible in both samples (Study 1: Unfamiliar = 85% [80/94], Not used = 13% [12/94], Used = 2% [2/94]; Study 2: Unfamiliar = 59% [41/69], Not used = 38% [26/69], Used = 3% [2/69]). Interestingly, trust towards algorithms in general was comparable in the two samples, and, on average, participants in both samples had neutral attitudes (measured on a scale of 0 to 10; Study 1: Mean (M) = 4.78, Standard Deviation (SD) = 2.18; Study 2: M = 5.20, SD = 2.00). Since our sample sizes were defined by resource constraints, we conducted sensitivity power analyses as per the recommended good practice Lakens (2022). The multilevel regressions directly comparing the SC with the OC or OT explanations (RQ2 and RQ3) were powered at the 80% level to detect effect sizes of Cohen‘s d of 0.31 and 0.38, in Study 1 and 2, respectively. [^9] We used GPower 3.1 for the sensitivity power analyses, and we calculated the effective sample sizes for each study by multiplying the number of participants by the cluster sizes (i.e. each user responded to two vignettes) and then dividing it by the design effect, which is 1+ (cluster size - 1) multiplied by the intra-class correlation (which we assumed to be low, r=0.2). Hence, we applied a design effect of 1.2 for our studies.
@@ 5 Discussion
Across two studies with different samples (from the general population and medical students), we found a reliable preference for Occlusion-1 over SHAP explanations. Our participants reported that they understood and trusted the former more than the latter, as expected. However, the comparison of all the explanations revealed that this effect may be solely driven by the fact that our participants found the OT explanations consistently more comprehensible than the SC explanations. Moreover, in Study 1, the OT explanations were also preferred to the SC explanations, though in Study 2 the results were less clear. These findings are consistent with a notion of people, or more specifically patients within a healthcare setting, preferring explanations in text over chart format Bertrand et al. (2023), for instance, due to finding some graphs difficult to interpret Durand et al. (2020); Szymanski et al. (2021). These findings indicate that prioritising FA explanations based on simple rather than complex models may not contribute to clarity. Interestingly, the results regarding trustworthiness were slightly in conflict with this idea, since in our medical student sample we only found evidence for higher trust in the OC compared to the SC explanations. While these findings show support for the role of explanation content, more research is needed to explore how people relate to and utilise different XAI techniques, and how and when an explanation’s format trumps the impact of its content. Future research should also explore the potential of tailoring explanation content and format to individuals Lesley and Hernández (2024), e.g. based on their expertise in a field Wang and Yin (2023) and their skills in interpreting numerical information (e.g. graph literacy Durand et al. (2020) or numeracy Reyna et al. (2009)).
@@
Our expert sample in Study 2 comprised medical students so it is not obvious how well our findings generalise to experienced healthcare practitioners: we leave this to future work. Also, self-report measures are prone to biases: Nagendran et al. Nagendran et al. (2023) found no correlation between self-reported understanding of XAI and the impact of XAI on behaviour, casting some doubt on their generalisability to real-life behaviour. Future research should investigate this presumed dissociation for behavioural measures of advice understanding and uptake Pálfi et al. (2022); Panigutti et al. (2022). This justifies our definition of a novel performance measure (i.e. the definition recognition task), especially since we found no evidence that participants were better at recognising the definition of either explanation: it would be interesting to study variants of this measure or different ways to assess it to better settle the question of the impact of explanation content. The task of choosing the correct definition of explanations may have been too difficult as performance was only reliably over chance level for the Occlusion-1 explanations in the medical student sample, potentially creating a floor effect, but, given the simple descriptions to choose from, this in itself calls into question user comprehension of FA explanations more generally.
@@
Ethical Statement
Participation in both studies was voluntary. Participants were asked to provide informed consent after reading an information sheet and they could withdraw at any time prior to completion, simply by closing the survey window, with their data deleted permanently. We informed the participants that once they had completed the survey, they could no longer withdraw and we would keep their data. All the participant responses were anonymised once payment was made (i.e. we deleted their e-mail addresses from the dataset). We received ethics approval from our institution for these studies.

### v3

@@ 4 Results
Study 1 was completed by 99 participants, from which 1 reported to have colour-blindness (we did not measure demographics). Since 5 participants failed the attention check, we did not include their data in any of our analyses (see more details below) resulting in a final sample of 94. Study 2 was completed by 69 medical students, from which 3 reported to have colour-blindness. [^10] One participant skipped the survey in Study 1 and one participant left a non-university email in Study 2; we deleted the data of both participants and did not include them in the results. The mean age of the participants in Study 2 was 22.3 (SD = 3.3), with 35 females (32 males, 1 other and 1 preferred not to disclose). As expected, participants in the medical student sample (Study 2) reported to have heard more about medical algorithms than participants in the Prolific sample (Study 1), but the number of people who had tried a medical AI tool before was negligible in both samples (Study 1: Unfamiliar = 85% [80/94], Not used = 13% [12/94], Used = 2% [2/94]; Study 2: Unfamiliar = 59% [41/69], Not used = 38% [26/69], Used = 3% [2/69]). Interestingly, trust towards algorithms in general was comparable in the two samples, and, on average, participants in both samples had neutral attitudes (measured on a scale of 0 to 10; Study 1: Mean (M) = 4.78, Standard Deviation (SD) = 2.18; Study 2: M = 5.20, SD = 2.00). Since our sample sizes were defined by resource constraints, we conducted sensitivity power analyses as per the recommended good practice [22]. The multilevel regressions directly comparing the SC with the OC or OT explanations (RQ2 and RQ3) were powered at the 80% level to detect effect sizes of Cohen‘s d of 0.31 and 0.38, in Study 1 and 2, respectively. [^11] We used GPower 3.1 for the sensitivity power analyses, and we calculated the effective sample sizes for each study by multiplying the number of participants by the cluster sizes (i.e. each user responded to two vignettes) and then dividing it by the design effect, which is 1+ (cluster size - 1) multiplied by the intra-class correlation (which we assumed to be low, r=0.2). Hence, we applied a design effect of 1.2 for our studies.
@@ 5 Discussion
Across two studies with different samples (from the general population and medical students), we found a reliable preference for Occlusion-1 over SHAP explanations. Our participants reported that they understood and trusted the former more than the latter, as expected. However, the comparison of all the explanations revealed that this effect may be solely driven by the fact that our participants found the OT explanations consistently more comprehensible than the SC explanations. Moreover, in Study 1, the OT explanations were also preferred to the SC explanations, though in Study 2 the results were less clear. These findings are consistent with a notion of people, or more specifically patients within a healthcare setting, preferring explanations in text over chart format [5], for instance, due to finding some graphs difficult to interpret [13, 40]. These findings indicate that prioritising FA explanations based on simple rather than complex models may not contribute to clarity. Interestingly, the results regarding trustworthiness were slightly in conflict with this idea, since in our medical student sample we only found evidence for higher trust in the OC compared to the SC explanations. While these findings show support for the role of explanation content, more research is needed to explore how people relate to and utilise different XAI techniques, and how and when an explanation’s format trumps the impact of its content. Future research should also explore the potential of tailoring explanation content and format to individuals [27], e.g. based on their expertise in a field [43] and their skills in interpreting numerical information (e.g. graph literacy [13] or numeracy [37]).
@@
Our expert sample in Study 2 comprised medical students so it is not obvious how well our findings generalise to experienced healthcare practitioners: we leave this to future work. Also, self-report measures are prone to biases: Nagendran et al. [31] found no correlation between self-reported understanding of XAI and the impact of XAI on behaviour, casting some doubt on their generalisability to real-life behaviour. Future research should investigate this presumed dissociation for behavioural measures of advice understanding and uptake [33, 35]. This justifies our definition of a novel performance measure (i.e. the definition recognition task), especially since we found no evidence that participants were better at recognising the definition of either explanation: it would be interesting to study variants of this measure or different ways to assess it to better settle the question of the impact of explanation content. The task of choosing the correct definition of explanations may have been too difficult as performance was only reliably over chance level for the Occlusion-1 explanations in the medical student sample, potentially creating a floor effect, but, given the simple descriptions to choose from, this in itself calls into question user comprehension of FA explanations more generally. Further, it would be interesting to study the effect of users’ awareness of XAI methods’ vulnerability to manipulation and capability for producing misleading explanations.